OpenStudy (anonymous):

If the line whose equation is y = x + 2k passes through point (3, -5) , then k = Select one: a. -4 b. 1 c. -2 d. 4

OpenStudy (anonymous):

a.-4

OpenStudy (anonymous):

because (x,y) =(3,-5) hence just put the values of 3 and 5 in it and you would get -5=3+2k -8=2k -4=k voila!